A cyclist accelerates from 0 m/s to 8 m/s in 3 seconds.What is his acceleration?Is this acceleration higher than that of a car which accelerates from 0 to 30 m/s in 8 seconds ?


1
Expert's answer
2022-03-18T09:15:15-0400

The acceleration of the car

"a=\frac{\Delta v}{\Delta t}"

Hence

"a_1=\frac{8\:\rm m/s}{3\:\rm s}=2.67\:\rm m/s^2"

"a_1=\frac{30\:\rm m/s}{8\:\rm s}=3.75\:\rm m/s^2"

"a_1<a_2"

Answer: No